# DEAR: Disentangled Environment and Agent Representations for Reinforcement Learning without Reconstruction

### v1

@@ I INTRODUCTION
Therefore, a key challenge for RL methods is to learn invariant representations that focus on the task-relevant aspects of the state and improve sample efficiency [6, 7, 8]. Previous methods proposed invariance based on reward through forward or inverse dynamics, i.e., two states are “behaviorally equivalent” when they yield identical immediate rewards and similar subsequent state distributions [6]; however, this approach does not perform well in challenging tasks [9]. Another strategy is to learn disentangled representations of observations, which separate the informative factors of variation in an image from their underlying causes [9, 2]. However, these methods do not explain how the disentangled representations are interpretable.
@@ III-A Disentangled Environment and Agent representations (DEAR)
Using the proposition in Equation. 1, the evidence lower bound on J can be expressed as:
@@ IV Experiments
Figure 5: Rewards for the three distracting DM control environments (left) Distracting Walker walk (middle) Distracting Cup catch (right) Distracting Finger spin. DEAR outperforms or matches other baselines on these tasks. Each curve is an average of 5 different seeds.
@@
    Robustness to Noise: Generally, the mask is noisy and approximate in real-world conditions since it is extracted from real sensor data, subject to noise and artifacts. Thus, we artificially introduce noise in the mask by generating patches of pixels around the joints. A noise coefficient of 0.8 is used, meaning 80% of noise is added. These masks were generated by randomly setting robot labels to non-robot labels with a user-specified probability. Additionally, an approximate mask with limited agent information is added. To generate an approximate mask, an original robot mask is downsampled and then upsampled back to its original size. Then, Gaussian blur and image thresholding are applied to get a new mask.
@@ V-A Benchmark control environment
Distracting Control Suite: The results indicate that DEAR performs strongly compared to SEAR in all three environments and obtains higher rewards, increasing the sample efficiency by at least 25%, shown in Fig. 5. This suggests that DEAR can successfully disentangle representations of the environment that are not relevant to control.
Figure 6: Training curves of different methods on two Franka kitchen tasks. 8 different seeds are used for each method.
@@ V-B Ablation study
Figure 7: Experiments with different $\beta$ values provide insights into the contributions of the objectives in the overall loss function and also the changes with different hyperparameters.
Robust to hyperparameters: We provide experiments to identify the optimal coefficients that influence the performance in Fig. 7. We found $\beta=100$ to be optimal for this task, and the results show some robustness for the values in the range $[1,100]$. Conversely, a negative coefficient leads to a reduced emphasis on disentanglement and the formation of overlapping vectors, which in turn diminishes performance. This indicates that weak disentanglement, characterized by highly overlapping vectors, hampers the agent’s ability to differentiate between the environment and itself, adversely affecting performance.
Robustness to noisy and approximate mask: We see, from Fig. 8, that the agent performance does not drop for both noisy and approximate masks, unlike SEAR, where a performance drop is reported [14]. This robustness will be the subject for future research activities, extending the tests to experiments with real robotic setups. Recent advancements in segmentation models like SAM [21] could be leveraged as a readily applicable solution to produce the segmentation mask.
Figure 8: (Top row) different variants of agent segmentation masks used. From left to right: Normal segmentation mask, noisy mask, and approximate mask. (Bottom row) training curves using noisy and approximate robot masks.
@@ VI CONCLUSIONS
One of the limitations of the current study is testing generalization to unseen environment parameters and adaptation to new tasks, which will be extended in future works. Moreover, we will investigate if spatiotemporal disentanglement could further optimize sample efficiency by providing a stronger understanding of the general physics of the environment and the agent.

### v2

@@ I INTRODUCTION
Therefore, a key challenge for RL methods is to learn invariant representations that focus on the task-relevant aspects of the state and improve sample efficiency i.e. learning to remove distractors from an image can provide a good prior for vision-based RL [6, 7, 8]. Previous methods proposed invariance based on reward through forward or inverse dynamics, i.e., two states are “behaviorally equivalent” when they yield identical immediate rewards and similar subsequent state distributions [6]; however, this approach does not perform well in challenging tasks [9]. Another strategy is to learn disentangled representations of observations, which separate the informative factors of variation in an image from their underlying causes [9, 2]. However, these methods do not explain how the disentangled representations are interpretable.
@@ III-A Disentangled Environment and Agent representations (DEAR)
Using the conjecture in Equation. 1, the evidence lower bound on J can be expressed as:
@@ IV Experiments
    Robustness to Noise: Generally, the mask is noisy and approximate in real-world conditions since it is extracted from real sensor data, subject to noise and artifacts. Thus, we artificially introduce noise in the mask by generating patches of pixels around the joints. A noise coefficient of 0.8 is used, meaning 80% of noise is added. These masks were generated by randomly setting robot labels to non-robot labels with a user-specified probability. Additionally, an approximate mask with limited agent information is added (See Fig. 5). To generate an approximate mask, an original robot mask is downsampled and then upsampled back to its original size. Then, Gaussian blur and image thresholding are applied to get a new mask.
    Figure 5: Different variants of agent segmentation masks used. From left to right: Normal segmentation mask, noisy mask, and approximate mask.
Figure 6: Training curves of different methods on three distracting DM control environments and two Franka kitchen tasks. DEAR outperforms or matches other baselines on these tasks. Each curve is an average of 5 different seeds.
@@ V-A Benchmark control environment
Distracting Control Suite: The results indicate that DEAR performs strongly compared to SEAR in all three environments and obtains a higher convergence for 1 million timesteps, increasing the sample efficiency by at least 25%, shown in Fig. 6. This suggests that DEAR can successfully disentangle representations of the environment that are not relevant to control.
@@ V-B Ablation study
Figure 7: (Left) Experiments with different $\beta$ values provide insights into the contributions of the objectives in the overall loss function and also the changes with different hyperparameters. Where $\alpha$ values are not provided, $\alpha=1$. Each experiment is an average of 5 random seeds. (Right) Training curves using noisy and approximate robot masks.
Robust to hyperparameters: We provide experiments to identify the optimal coefficients that influence the performance in Fig. 7 (left). We found $\beta=100$ to be optimal for this task, and the results show some robustness for the values in the range $[1,100]$. Conversely, a negative coefficient leads to a reduced emphasis on disentanglement and the formation of overlapping vectors, which in turn diminishes performance. This indicates that weak disentanglement, characterized by highly overlapping vectors, hampers the agent’s ability to differentiate between the environment and itself, adversely affecting performance.
Robustness to noisy and approximate mask:
We see, from Fig. 7(Right), that the agent performance does not drop for both noisy and approximate masks, unlike SEAR, where a performance drop is reported [14]. This robustness will be the subject for future research activities, extending the tests to experiments with real robotic setups. Recent advancements in segmentation models like SAM [21] could be leveraged as a readily applicable solution to produce the segmentation mask.
@@ VI CONCLUSIONS
One of the limitations of the current study is testing generalization to unseen environment parameters and adaptation to new tasks such as dynamic backgrounds, which will be extended in future works. Moreover, we will investigate if spatiotemporal disentanglement could further optimize sample efficiency by providing a stronger understanding of the general physics of the environment and the agent.
For real-world applicability, some new segmentation models such as SAM has shown promising performance on out-of-distribution robot/agent segmentation masks [21]. Using such models to generate the segmentation masks for supervision can show robust performance in real-world environments that are more noisier and dynamic than the controlled simulation settings. This translation from simulator to real would robotic agents will be part of upcoming works.